OpenStudy (anonymous):

x^3.5=46.74 solve of x

OpenStudy (campbell_st):

use logs ln(x^3.5) = ln(46.74) applying log laws \[3.5\ln(x) = \ln(46.74)\] divide both sides by 3.5 evaluate the right hand side the raise each side as a power of e for the solution.

OpenStudy (anonymous):

\[2^{2} = 4...............\sqrt{4} = 2\] same way \[\sqrt[3.5]{46.74} = 2.99\] so x = 2.99

OpenStudy (anonymous):

how do you type that into a TI-84 calculator?

OpenStudy (anonymous):

the 3.5 part

OpenStudy (anonymous):

hmm... dont have a TI-84 calc.. but in mine there is a filled squarebox beside a square root, which has a unfilled squarebox

OpenStudy (anonymous):

ohh hmm I don't have that but I figured it out! Thank you both!

OpenStudy (anonymous):

npz